### v1

# Formations of generalized Wannier-Stark ladders: Theorem and applications

Abstract
The Wannier-Stark ladder (WSL) is a basic concept, supporting periodic oscillation, widely used in many areas of physics. In this paper, we investigate the formations of WSL in generalized systems, including strongly correlated and non-Hermitian systems. We present a theorem on the existence of WSL for a set of general systems that are translationally symmetric before the addition of a linear potential. For a non-Hermitian system, the WSL becomes complex but maintains a real energy level spacing. We illustrate the theorem using 1D extended Bose-Hubbard models with both real and imaginary hopping strengths. It is shown that the Bloch-Zener oscillations of correlated bosons are particularly remarkable under resonant conditions. Numerical simulations for cases with boson numbers $n=2$, $3$, and $4$ are presented. Analytical and numerical results for the time evolution of the $n$-boson-occupied initial state indicate that all evolved states exhibit quasi periodic oscillations, but with different profiles, depending on the Hermiticity and interaction strength.
I Introduction
Rigorous results in quantum many-body systems are rare but crucial for gaining valuable insights into the characterization of dynamic behaviors within correlated systems. For non-interacting system, Wannier-Stark (WS) ladders are a theoretical model in solid-state physics that describe the quantization of electronic states in a crystal under the influence of a constant electric field Bloch (1929); Wannier (1959, 1960); Glück (2002); Waschke et al. (1993). This has been observed in systems of semiconductor superlattices and ultracold atoms Waschke et al. (1993); Ben Dahan et al. (1996); Wilkinson et al. (1996); Anderson and Kasevich (1998). It is closely related to Bloch oscillations (BOs) due to the equidistant spectrum. In the past decade, such a phenomenon has attracted much attention in cold-atoms physics and photonics due to applications in interferometric measurements and as a method for manipulating localized wave packets Breid et al. (2006, 2007); Dreisow et al. (2009); Kling et al. (2010); Plötz and Wimberger (2011). It can be simulated using artificial quantum systems, such as superconducting circuits Song et al. (2024). On the other hand, the dynamics of particle pairs in lattice systems have garnered considerable interest, owing to the rapid advancements in experimental techniques. Ultra-cold atoms have proven to be an ideal testing ground for few-particle fundamental physics, as optical lattices offer clean realizations of a variety of many-body Hamiltonians. It stimulates many experimental Winkler et al. (2006); Fölling et al. (2007); Gustavsson et al. (2008) and theoretical investigations Mahajan and Thyagaraja (2006); Petrosyan et al. (2007); Creffield (2007); Kuklov and Moritz (2007); Zöllner et al. (2008); Wang et al. (2008); Valiente and Petrosyan (2008); Jin et al. (2009); Valiente and Petrosyan (2009); Valiente et al. (2010); Javanainen et al. (2010); Wang and Liang (2010); Rosch et al. (2008); Zhang (2024) in strongly correlated systems. Recently, it has been shown that a bound pair within a correlated system also exhibits periodic dynamics when subjected to a linear potential Khomeriki et al. (2010); Longhi (2011, 2012); Corrielli et al. (2013); Lin et al. (2014); Zhang et al. (2016); Zhang (2024).
In this work, we extend the concept of WSL to the many-body systems. We investigate the formations of WSL in generalized systems, including strongly correlated and non-Hermitian systems. We present a theorem on the existence of WSL for a set of general systems that are translationally symmetric before the addition of a linear potential. We show that the WSL becomes complex but maintains a real energy level spacing for a non-Hermitian system. To illustrate the theorem, we focus on 1D extended Bose-Hubbard models with both real and imaginary hopping strengths. It is shown that the Bloch-Zener oscillations of correlated bosons are particularly remarkable under resonant conditions. Numerical simulations for cases with boson numbers $n=2$, $3$, and $4$ are presented. Analytical and numerical results for the time evolution of the $n$-boson-occupied initial state indicate that all evolved states exhibit quasi periodic oscillations, but with different profiles, depending on the Hermiticity and interaction strength.
This paper is organized as follows. In Sec. II, we present the theorem on the structure of energy levels for general systems in a linear field, including Hermitian and non-Hermitian. In Sec. III, we take the extended Bose-Hubbard model as an application of the theorem and provide its effective Hamiltonian. Sec. IV and Sec. V are dedicated to the dynamics of the multi-particle effective Hamiltonian and original Hamiltonian, respectively. Finally, we provide a summary in Sec. VI.
II Theorem on WSLs
In general, a linear field can break the translational symmetry and disrupt the energy band structure of a non-interacting system. For some simple systems, it has been precisely demonstrated that the formation of WSLs leads to periodic dynamics, specifically Bloch oscillations. However, obtaining the energy levels for complex systems, particularly those with interactions, is quite challenging. The main aim of this paper is to determine whether WSLs still exist in general systems. In this section, we present a theorem on the structure of energy levels for general systems. We focus solely on the 1D system for simplicity. The obtained conclusion is applicable to higher-dimensional systems.
We start with a general tight-binding model with the Hamiltonian in the form
$H=H_{0}+\omega\sum_{j=-\infty}^{\infty}\sum_{\sigma=1}^{\Lambda}ja_{j,\sigma}^{{\dagger}}a_{j,\sigma},$ (1)
where $a_{j,\sigma}^{{\dagger}}$ ($a_{j,\sigma}$) is the boson or fermion creation (annihilation) operator, with internal degree of freedom (or flavor) $\sigma=\left[1,\Lambda\right]$, at the $i$th site.
Theorem.—The energy levels of Hamiltonian $H$ must consist of multi-set of WSLs with an identical real level spacing, which is independent of $H_{0}$, if $H_{0}$ is constructed under the following conditions:
(i) The total particle number $\sum_{j=-\infty}^{\infty}\sum_{\sigma=1}^{\Lambda}a_{j,\sigma}^{{\dagger}}a_{j,\sigma}$ is conservative, that is
$[\sum_{j=-\infty}^{\infty}\sum_{\sigma=1}^{\Lambda}a_{j,\sigma}^{{\dagger}}a_{j,\sigma},H_{0}]=0.$ (2)
(ii) Hamiltonian $H_{0}$ has translational symmetry, that is
$\left[T_{n_{0}},H_{0}\right]=0,$ (3)
where $T_{n_{0}}$ is the translational operator defined as
$T_{n_{0}}a_{j,\sigma}T_{n_{0}}^{-1}=a_{j+n_{0},\sigma}.$ (4)
Proof.–Based on the assumptions on $H$, we simply have
$\left[T_{r_{0}},H\right]=-r_{0}\omega\sum_{j=-\infty}^{\infty}\sum_{\sigma=1}^{\Lambda}a_{j,\sigma}^{{\dagger}}a_{j,\sigma},$ (5)
which is referred to as a ramped translational symmetry. in the invariant subspace with fixed particle number $n_{0}$, any eigenstate of $H$ obeys
$H|\psi_{0}\rangle=E_{0}|\psi_{0}\rangle,$ (6)
and
$\sum_{j=-\infty}^{\infty}\sum_{\sigma=1}^{\Lambda}a_{j,\sigma}^{{\dagger}}a_{j,\sigma}|\psi_{0}\rangle=n_{0}|\psi_{0}\rangle.$ (7)
Then we always have
$H\left(T_{r_{0}}|\psi_{0}\rangle\right)=\left(E_{0}+r_{0}\mathcal{\omega}\right)\left(T_{r_{0}}|\psi_{0}\rangle\right),$ (8)
and
$H\left(T_{r_{0}}^{-1}|\psi_{0}\rangle\right)=\left(E_{0}-r_{0}\mathcal{\omega}\right)\left(T_{r_{0}}|\psi_{0}\rangle\right),$ (9)
i.e., $T_{r_{0}}|\psi_{0}\rangle$ ($T_{r_{0}}^{-1}|\psi_{0}\rangle$) is also the eigenstate of $H$ with eigen energy $E_{0}+r_{0}\mathcal{\omega}$ ($E_{0}-r_{0}\mathcal{\omega}$). Operator $T_{r_{0}}$ increments the energy level by a step of $r_{0}\mathcal{\omega}$ and the operator $T_{r_{0}}^{-1}$ decrements the energy level by a step of $r_{0}\mathcal{\omega}$. We can then construct a set of eigenstates
$|\psi_{m}\rangle=\left(T_{r_{0}}\right)^{m}|\psi_{0}\rangle,$ (10)
$(m=0,\pm 1,\pm 2,...)$ with eigenenergy
$E_{m}=E_{0}+mr_{0}\mathcal{\omega}.$ (11)
It indicates that the translational operator $T_{r_{0}}$ acts as a ladder operator. This proof is independent of $H_{0}$. We note that based on another eigenstate $|\psi_{0}^{\prime}\rangle$, which does not be long to $\left\{|\psi_{m}\rangle\right\}$, another set of eigenstates $\left\{|\psi_{m}^{\prime}\rangle\right\}$ can be generated accordingly. Although $E_{0}$ can be real or complex, dependenting on the detail of the Hamiltonian $H_{0}$, we can conclude that the energy levels of Hamiltonian $H$ consists of multi-set of WSLs with an identical real level spacing $r_{0}\mathcal{\omega}$.
Applying the conclusion to the several cases, the same result can be obtained in an exact manner. (i) Considering the fermonic Hubbard model, where the flavor $\Lambda=2$, with $a_{j,1}=c_{j,\uparrow}$ and $a_{j,2}=c_{j,\downarrow}$ being the fermion operators, the doublon Bloch oscillation is the direct demonstration of the existence of WSLs Longhi (2012); Zhang (2024). (ii) Considering the extended Bose-Hubbard model, where the flavor $\Lambda=1$, with $a_{j,1}=b_{j}$ being boson operator, the bound pair Bloch oscillation is also the direct demonstration of the existence of WSLs Lin et al. (2014); Zhang et al. (2016).
This rigorous conclusion has important implications for investigating the dynamics of many-body and non-Hermitian systems in the presence of a linear field. In the following sections, we will focus on both the Hermitian and non-Hermitian extended Bose-Hubbard models to demonstrate the application of our conclusions and to reveal the dynamics of correlated bosons.
III Extended Bose-Hubbard model
We consider an extended boson Hubbard model describing interacting particles in the lowest Bloch band of a one dimensional lattice, which can be employed to describe ultracold atoms or molecules with magnetic or electric dipole-dipole interactions in optical lattices. We focus on the dynamics of the bosonic cluster, which is $n$ identical bosons in a bound state. For the simplest case with $n=2$, it has been demonstrated that, as another type of bound pair, it allows for correlated single-particle tunneling, as shown in previous work Jin and Song (2011); Lin et al. (2014). Such a bound pair can act as a quasi-particle, with an energy band width of the same order as that of a single particle. One of the objectives of this paper is to demonstrate that this type of bound state can be generalized to an $n$-boson system.
We consider an illstrative example that $H_{0}$ is a one-dimensional extended Bose-Hubbard model with the Hamiltonian
$H_{\text{{B}}}$ $=$ $H_{0}+\omega\sum_{j=-\infty}^{\infty}jn_{j},$
[…]
$+\sum_{j=-\infty}^{\infty}[\frac{U}{2}n_{j}\left(n_{j}-1\right)+Vn_{j}n_{j+1}],$
where $b_{j}^{{\dagger}}$ and $n_{j}=b_{j}^{{\dagger}}b_{j}$ is the creation and number operators of the boson at the $j$th site. Parameters $\kappa$, $U$ and $V$, denote the tunneling strength, on-site and NN interactions between bosons. In this work, we consider the cases with real and imaginary , respectively. Obviously, Hamiltonian $H_{0}$ satisfies the conditions in the theorem, and then possesses multi-set of WSLs.
Although we cannot obtain the value of $E_{0}$ for each set of WSL, additional some features of the energy levels can be revealed for the systems under some conditions. In the following, we will consider two typical cases.
III.1 $\mathcal{PT}$ symmetry
We introduce two operators. One is a linear operator $\mathcal{P}$, which represents the local gauge transformation
$\mathcal{P}b_{l}\mathcal{P}^{-1}=(-1)^{l}b_{l}.$ (13)
The other is an anti-linear operator $\mathcal{T}$, which is the time-reversal operator defined as
$\mathcal{T}\sqrt{-1}\mathcal{T}^{-1}=-\sqrt{-1}.$ (14)
We note that the extended Bose-Hubbard $H_{\text{{B}}}$ has $\mathcal{PT}$ symmetry, that is
$\mathcal{PT}H_{\text{{B}}}\mathcal{(PT)}^{-1}=H_{\text{{B}}},$ (15)
when we consider the imaginary hopping strength, $\kappa=i\left|\kappa\right|$. According to non-Hermitian quantum mechanics Scholtz et al. (1992), the Hamiltonian $H_{\text{{B}}}$ is pseudo-Hermitian, in which complex energy levels appear in the conjugate pairs. This provides a constraint on the complex WSLs in the Hamiltonian $H_{\text{{B}}}$, which will be utilized in the investigation of dynamics in the following section.
III.2 Resonant interactions
Now we consider the Hamiltonian $H_{\text{{B}}}$ under the resonant condition $U=V\gg\left|\kappa\right|$. Let us begin by analyzing in detail the many-boson problem within the Hamiltonian $H_{0}$ with zero $\kappa$. It is clear that, in the invariant subspace with fixed boson number $n$, there is a set of degenerate eigenstates $\left\{\left|l\right\rangle,l\in\left(-n\infty,n\infty\right)\right\}$ of $H_{0}\left(\kappa=0\right)$ with eigen energy $\varepsilon(n)$ $=Un(n-1)/2$, which are expressed as the form
$\left|jn+\lambda\right\rangle=\frac{\left(b_{j}^{\dagger}\right)^{n-\lambda}\left(b_{j+1}^{\dagger}\right)^{\lambda}}{\sqrt{\left(n+1-\lambda\right)!\left(\lambda-1\right)!}}|\mathrm{vac}\rangle,$ (16)
with $j\in\left(-\infty,\infty\right)$ and $\lambda\in[0,n-1]$. Here $\left|\mathrm{vac}\right\rangle$ is the vacuum state for the boson operator $b_{j}$. It is easy to check that the nearest energy level next to $\varepsilon(n)$ is $\varepsilon(n)-U$, resulting in an energy gap of $U$. This ensures us to perform the first-order perturbation approximation when the hopping term is considered under the strong correlation condition where $U\gg\left|\kappa\right|$.
In order to obtain the effective Hamiltonian, one can simply envalue the nonzero matrix elements of $H_{\text{{B}}}$ in the subspace spanned by the set of eigenstates $\left\{\left|l\right\rangle\right\}$. For a given boson number $n$, the effective Hamiltonian can be written as
$H_{\mathrm{eff}}^{[n]}$ $=$ $-\kappa\sum_{j,\lambda}h_{j,\lambda}+U\frac{n(n-1)}{2}\sum_{l=-\infty}^{\infty}\left|l\right\rangle\left\langle l\right|$ (17)
$+\omega\sum_{l=-\infty}^{\infty}l\left|l\right\rangle\left\langle l\right|,$
where
$h_{j,\lambda}$ $=$ $\sqrt{(n-\lambda)(\lambda+1)}\left|jn+\lambda\right\rangle$ (18)
[…]
In its present form, $H_{\mathrm{eff}}^{[n]}$ are formally analogous to a tight-binding model describing a single-particle dynamics in an infinite chain with NN hopping, under a linear field. We note that the states $\left|l\right\rangle$ in each unit cell represent an $n$-bosonic-cluster state, in which all $n$ bosons occupy a single dimer of the original lattice of the Bose-Hubbard system $H_{\text{{B}}}$. In this sense, the dynamics of $H_{\mathrm{eff}}^{[n]}$ is essentially the dynamics of $n$ correlated bosons. Notably, we would like to point out that the hopping strength in $H_{\mathrm{eff}}^{[n]}$ is of the order of $\kappa$, not $\kappa^{2}/U$. This ensures that the phenomena arising from the effective Hamiltonian can be observed in experiments, similar to those of a single boson. On the other hand, our theorem also holds for the effective Hamiltonian $H_{\mathrm{eff}}^{[n]}$, which describes a single-particle system of $n$-site unit cell. The corresponding step for each ladder is $n\omega$.
Figure 1: Plots of $p_{l}(t)$ defined in Eq. (24) and $P_{j}(t)$ defined in Eq. (27), obtained by numerical diagonalizations. Here, (a1) and (a2) correspond to the effective Hamiltonian $H_{\mathrm{eff}}^{[3]}$ with $\kappa=1.0$ and $1.0i$, respectively. (b1) and (b2) correspond to the Hamiltonian $H_{\mathrm{B}}$ with $\kappa=1.0$ and $1.0i$, respectively. Other system paraneters are $\omega=0.1$, $U=V=5.0$, and $n=3$. It indicates that $H_{\mathrm{eff}}^{[3]}$ and $H_{\mathrm{B}}$ have the similar dynamics, demonstrating a multi-ladder energy level structure.
Figure 2: The same plots as Fig. 1 for the case with $n=4$.
Figure 3: Plots of the dressed energy $\varepsilon$ as a function of $\kappa$, given by Eq. (25) for understanding the energy level structures of the extended Bose-Hubbard model subjected to a linear potential. Several representative energy levels for the systems are selected: (a) $n=3$ and (b) $n=4$. The system parameters are $U=V=5.0$ and $\omega=0.1$. The black dashed lines are added to indicate the relations between the curves. The orange dashed lines are added to indicate the coalescing WSL. We see that the patterns consist of three and four families of curves, respectively.
IV Dynamics of effective Hamiltonian $H_{\mathrm{eff}}^{[n]}$
In this section, we demonstrate the application of the theorem to the effective Hamiltonian $H_{\mathrm{eff}}^{[n]}$ with small $n$. We will present the explicit form of $H_{\mathrm{eff}}^{[n]}$, analyze the relative energy structures among the sets of ladders based on the symmetry of the system, and then perform numerical simulations to study the corresponding dynamics.
Introducing a reflection operator $\mathcal{R}$, which is a unitary operator and defined as
$\mathcal{R}\left|l\right\rangle=(-1)^{l}\left|-l\right\rangle$ (19)
we have
$\mathcal{R}H_{\mathrm{eff}}^{[n]}\mathcal{R}^{-1}=-H_{\mathrm{eff}}^{[n]}.$ (20)
It indicates that the spectrum is symmetric about zero energy because a unitary transformation preserves the spectrum of the system. This conclusion holds for the system with an arbitrary value of $\kappa$, including imaginary numbers, which is associated with complex energy levels. In addition, $H_{\mathrm{eff}}^{[n]}$ still possesses $\mathcal{PT}$ symmetry, given by the relation $\mathcal{PT}H_{\mathrm{eff}}^{[n]}\mathcal{(PT)}^{-1}=H_{\mathrm{eff}}^{[n]}$. This ensures that Hamiltonian $H_{\mathrm{eff}}^{[n]}$ is pseudo-Hermitian, meaning that complex energy levels appear in the conjugate pairs, according to non-Hermitian quantum mechanics Scholtz et al. (1992). Therefore, the energy levels of $H_{\mathrm{eff}}^{[n]}$ must satisfy the aforementioned constraints.
In general, the number of the sets of energy ladders is equal to $n$, or the site number of the unit cell. This feature should be clearly demonstrated in the dynamics by the occurrence of normal Bloch oscillations and Bloch-Zener oscillations. In the following, we will detail the level structures for both the Hermitian and non-Hermitian Hamiltonians $H_{\mathrm{eff}}^{[n]}$, focusing on small values of $n$.
(i) In the case where $n=2$, the effective Hamiltonian is given by
$H_{\mathrm{eff}}^{[2]}=-\sqrt{2}\kappa\sum_{l=-\infty}^{\infty}(|l\rangle\langle l+1|+\mathrm{H.c.})+\omega\sum_{l=-\infty}^{\infty}l\left|l\right\rangle\left\langle l\right|.$ (21)
This represents the simplest WSL system, which emerges from a uniform chain possessing a single real energy ladder with spacing $\omega$, regardingless of whether $\kappa$ is real or imaginary.
(ii) In the case where $n=3$, the effective Hamiltonian is given by
$H_{\mathrm{eff}}^{[3]}$ $=$ $-\kappa\sum_{j=-\infty}^{\infty}(\sqrt{3}|3j\rangle\langle 3j+1|+2|3j+1\rangle\langle 3j+2|$ (22)
$+\sqrt{3}|3j+2\rangle\langle 3j+3|+\mathrm{H.c.})$
$+\omega\sum_{l=-\infty}^{\infty}l\left|l\right\rangle\left\langle l\right|.$
There are three sets of real ladders with spacing $3\omega$ when $\kappa$ is real. One of them contains a zero energy level, while all the levels of other two ladders are symmetric about the zero energy.
On the other hand, when $\kappa$ is imaginary, there are still three sets of ladders with a spacing $3\omega$. The aforementioned constraints necessitate the existence of zero energy. Numerical results show that the other two are complex conjugates of each other.
(iii) In the case where $n=4$, the Hamiltonian is given by
$H_{\mathrm{eff}}^{[4]}$ $=$ $-\kappa\sum_{j=-\infty}^{\infty}(2|4j\rangle\langle 4j+1|+\sqrt{6}|4j+1\rangle\langle 4j+2|\,$ (23)
$+\sqrt{6}|4j+2\rangle\langle 4j+3|\,+2|4j+3\rangle\langle 4j+4|+\mathrm{H.c.})$
$+\omega\sum_{l=-\infty}^{\infty}l\left|l\right\rangle\left\langle l\right|.$
Numerical simulation shows that, when $\kappa$ is imaginary, there are two sets of real ladders and two sets of complex ladders. The two real ladders form a new ladder with energy levels given by ±$\pm 2\omega m$, where $m=0$, $1$, …. The real parts of two complex ladders is given by $(2\pm 4m)\omega$, where $m=0$, $1$, ….
These results lead to the conclusion that for small $n$, in the case with imaginary $\kappa$, there exists at least one set of real energy ladder. This suggests that normal Bloch oscillations can be observed in such non-Hermitian systems for a carefully chosen initial state. In contrast, an arbitrary initial state should not exhibit a regular Bloch oscillation due to the imaginary parts of the energy levels. However, the component of the evolved state that involves the ladder with the maximal imaginary part should dominate after a period of time. To demonstrate these points, numerical simulations are performed for the dynamics driven by two typical effective Hamiltonians. We consider the time evolution of an initial state $\left|\Psi\left(0\right)\right\rangle$, which is a site state corresponding to the multi-boson state, given by $\frac{1}{\sqrt{n!}}\left(b_{j}^{\dagger}\right)^{n}|\mathrm{vac}\rangle$. To characterize the profile of the evolved state, we compute the Dirac probability distribution, defined as
$p_{l}(t)=\left|\left\langle l\right|\Psi\left(t\right)\rangle\right|^{2}=\left|\left\langle l\right|e^{-i\Lambda_{\mathrm{\max}}^{[n]}t}e^{-iH_{\mathrm{eff}}^{[n]}t}\left|\Psi\left(0\right)\right\rangle\right|^{2}.$ (24)
[…]
Before we conclude this section, we would like to point out that the aforementioned results are obtained from effective Hamiltonians, which describe non-interacting systems and subspaces of correlated bosons in the large $U$ limit.
V Dynamics of correlated bosons
In this section, we focus on the investigation on the dynamics of correlated bosons. Considering the original Bose-Hubbard Hamiltonian as given by Eq. (12), the theorem regarding its energy levels applies to any set of parameters $\kappa$, $U$ and $V$. This feature can be characterized by the quantity $\varepsilon$, which is defined as the sum of the real part of $E$ and the imaginary part of $E$,
[…]
This quantity is referred to as dressed energy. Although $\varepsilon$ does not have an evident physical meaning, it is still expressed in what is known as the ladder form
$\varepsilon=\varepsilon_{0}\pm m\omega,$ (26)
where $m=0$, $1$, $2$, …. The advantage of $\varepsilon$ is that one can demonsrate the complex ladder using the real ladder of $\varepsilon$. When an exceptional point appears in the complex region of $E$, we can directly observe the coalescence of two sets of $\varepsilon$ ladder. It is remarkable that the $\varepsilon$ ladders also exhibit level repulsion around the exceptional points. To demonstrate these features we compute the quantity $\varepsilon$ as a function of $\kappa$, which is obtained by the exact diagonalization of the Hamiltonian $H_{\mathrm{B}}$ on finite size chains with $n=3$ and $4$. In Fig. 3, quantities $\varepsilon(\kappa)$ are plotted for both real and imaginary $\kappa$. As expected, the plots of $\varepsilon(\kappa)$ are evidently several families of curves. A coalescing ladder appears in the imaginary region of $\kappa$.
Based on these observations, one can detect the existence of such ladder structure by the dynamical behavior of a given state. To demonstrate these points, numerical simulations are performed for the dynamics driven by the original Hamiltonian. We consider the time evolution of an initial state $\left|\Psi\left(0\right)\right\rangle=\frac{1}{\sqrt{n!}}\left(b_{j}^{\dagger}\right)^{n}|\mathrm{vac}\rangle$, which is multi-boson state. To characterize the profile of the evolved state, we compute the Dirac probability distribution, defined as
$P_{j}(t)=\left\langle\Psi\left(t\right)\right|a_{j}^{{\dagger}}a_{j}\left|\Psi\left(t\right)\right\rangle=\left|a_{j}e^{-iH_{\text{B}}t}e^{-\Lambda_{\mathrm{\max}}t}\left|\Psi\left(0\right)\right\rangle\right|^{2}$ (27)
[…]
As predicted, we observe (i) Bloch-Zener oscillations for Hermitian systems; (ii) stable oscillations after a period of time for non-Hermitian systems. This indicates that the dynamics for the Hamiltonian $H_{\mathrm{B}}$ with moderate $U$ is similar to that for the effective Hamiltonian $H_{\mathrm{eff}}^{[n]}$. Such behavior provides a dynamic signature indicative of the existence of the WSLs in the original Bose-Hubbard model.
VI Summary
In summary, we have developed a theory for a class of Hermitian and non-Hermitian Hamiltonians that possess WSLs. Such Hamiltonians consist of two distinct parts: an arbitrary Hamiltonian exhibiting translational symmetry and a linear potential. The most fascinating and important feature of these systems is that their conclusions are independent of the details of the Hamiltonian. As examples, we have investigate the extended Bose-Hubbard model with several bosons. Analytic and numerical approaches verified our conclusion in the case with $U=V$. We also demonstrated the application of the theorem by analyzing dynamics driven by the Hubbard model using numerical simulations. Our findings provide a method for analyzing the energy-level structure of general systems and are expected to be both necessary and insightful for quantum engineering in many-body systems.
Acknowledgements.
This work was supported by National Natural Science Foundation of China (under Grant No. 12374461).

### v2

@@
 I Introduction
-Rigorous results in quantum many-body systems are rare but crucial for gaining valuable insights into the characterization of dynamic behaviors within correlated systems.
-For non-interacting system, Wannier-Stark (WS) ladders are a theoretical model in solid-state physics that describe the quantization of electronic states in a crystal under the influence of a constant electric field Bloch (1929); Wannier (1959, 1960); Glück (2002); Waschke et al.
+Rigorous results in quantum interacting systems are rare but crucial for gaining valuable insights into the characterization of dynamic behaviors within correlated systems.
+For non-interacting system, Wannier-Stark ladders (WSL) are a theoretical model in solid-state physics that describe the quantization of electronic states in a crystal under the influence of a constant electric field Bloch (1929); Wannier (1959, 1960); Glück (2002); Waschke et al.
 (1993).
@@
 On the other hand, the dynamics of particle pairs in lattice systems have garnered considerable interest, owing to the rapid advancements in experimental techniques.
-Ultra-cold atoms have proven to be an ideal testing ground for few-particle fundamental physics, as optical lattices offer clean realizations of a variety of many-body Hamiltonians.
+Ultra-cold atoms have proven to be an ideal testing ground for few-particle fundamental physics, as optical lattices offer clean realizations of a variety of interacting Hamiltonian systems.
 It stimulates many experimental Winkler et al.
@@
 (2016); Zhang (2024).
-In this work, we extend the concept of WSL to the many-body systems.
+In this work, we extend the concept of WSL to the interacting Hamiltonian systems.
 We investigate the formations of WSL in generalized systems, including strongly correlated and non-Hermitian systems.
@@
 In Sec.
-III, we take the extended Bose-Hubbard model as an application of the theorem and provide its effective Hamiltonian.
+III, we take the extended Bose-Hubbard model as an application of the theorem and provide its effective Hamiltonian under the resonant condition in strongly interaction limit.
 Sec.
-IV and Sec.
-V are dedicated to the dynamics of the multi-particle effective Hamiltonian and original Hamiltonian, respectively.
+IV dedicated to the dynamics of the few-boson effective Hamiltonian and the corresponding original Hamiltonian, respectively.
+In particular, we also provide numerical simulations for both Hermitian and non-Hermitian sample systems, which offer evident demonstrations of our rigorous results, in Sec.
+V.
 Finally, we provide a summary in Sec.
@@
 (ii) Hamiltonian $H_{0}$ has translational symmetry, that is
-$\left[T_{n_{0}},H_{0}\right]=0,$ (3)
-where $T_{n_{0}}$ is the translational operator defined as
-$T_{n_{0}}a_{j,\sigma}T_{n_{0}}^{-1}=a_{j+n_{0},\sigma}.$ (4)
+$\left[T_{r},H_{0}\right]=0,$ (3)
+where $T_{r}$ is the translational operator defined as
+$T_{r}a_{j,\sigma}T_{r}^{-1}=a_{j+r,\sigma}.$ (4)
 Proof.–Based on the assumptions on $H$, we simply have
-$\left[T_{r_{0}},H\right]=-r_{0}\omega\sum_{j=-\infty}^{\infty}\sum_{\sigma=1}^{\Lambda}a_{j,\sigma}^{{\dagger}}a_{j,\sigma},$ (5)
-which is referred to as a ramped translational symmetry. in the invariant subspace with fixed particle number $n_{0}$, any eigenstate of $H$ obeys
+$\left[T_{r},H\right]=-r\omega\sum_{j=-\infty}^{\infty}\sum_{\sigma=1}^{\Lambda}a_{j,\sigma}^{{\dagger}}a_{j,\sigma}T_{r},$ (5)
+which is referred to as a ramped translational symmetry. in the invariant subspace with fixed particle number $N$, any eigenstate of $H$ obeys
 $H|\psi_{0}\rangle=E_{0}|\psi_{0}\rangle,$ (6)
 and
-$\sum_{j=-\infty}^{\infty}\sum_{\sigma=1}^{\Lambda}a_{j,\sigma}^{{\dagger}}a_{j,\sigma}|\psi_{0}\rangle=n_{0}|\psi_{0}\rangle.$ (7)
+$\sum_{j=-\infty}^{\infty}\sum_{\sigma=1}^{\Lambda}a_{j,\sigma}^{{\dagger}}a_{j,\sigma}|\psi_{0}\rangle=N|\psi_{0}\rangle.$ (7)
 Then we always have
-$H\left(T_{r_{0}}|\psi_{0}\rangle\right)=\left(E_{0}+r_{0}\mathcal{\omega}\right)\left(T_{r_{0}}|\psi_{0}\rangle\right),$ (8)
+$H\left(T_{r}|\psi_{0}\rangle\right)=\left(E_{0}+Nr\mathcal{\omega}\right)\left(T_{r}|\psi_{0}\rangle\right),$ (8)
 and
-$H\left(T_{r_{0}}^{-1}|\psi_{0}\rangle\right)=\left(E_{0}-r_{0}\mathcal{\omega}\right)\left(T_{r_{0}}|\psi_{0}\rangle\right),$ (9)
-i.e., $T_{r_{0}}|\psi_{0}\rangle$ ($T_{r_{0}}^{-1}|\psi_{0}\rangle$) is also the eigenstate of $H$ with eigen energy $E_{0}+r_{0}\mathcal{\omega}$ ($E_{0}-r_{0}\mathcal{\omega}$).
-Operator $T_{r_{0}}$ increments the energy level by a step of $r_{0}\mathcal{\omega}$ and the operator $T_{r_{0}}^{-1}$ decrements the energy level by a step of $r_{0}\mathcal{\omega}$.
+$H\left(T_{r}^{-1}|\psi_{0}\rangle\right)=\left(E_{0}-Nr\mathcal{\omega}\right)\left(T_{r}^{-1}|\psi_{0}\rangle\right),$ (9)
+i.e., $T_{r}|\psi_{0}\rangle$ ($T_{r}^{-1}|\psi_{0}\rangle$) is also the eigenstate of $H$ with eigen energy $E_{0}+Nr\mathcal{\omega}$ ($E_{0}-Nr\mathcal{\omega}$).
+Operator $T_{r}$ increments the energy level by a step of $Nr\mathcal{\omega}$ and the operator $T_{r}^{-1}$ decrements the energy level by a step of $Nr\mathcal{\omega}$.
 We can then construct a set of eigenstates
-$|\psi_{m}\rangle=\left(T_{r_{0}}\right)^{m}|\psi_{0}\rangle,$ (10)
+$|\psi_{m}\rangle=\left(T_{r}\right)^{m}|\psi_{0}\rangle,$ (10)
 $(m=0,\pm 1,\pm 2,...)$ with eigenenergy
-$E_{m}=E_{0}+mr_{0}\mathcal{\omega}.$ (11)
-It indicates that the translational operator $T_{r_{0}}$ acts as a ladder operator.
+$E_{m}=E_{0}+mNr\mathcal{\omega}.$ (11)
+It indicates that the translational operator $T_{r}$ acts as a ladder operator.
 This proof is independent of $H_{0}$.
 We note that based on another eigenstate $|\psi_{0}^{\prime}\rangle$, which does not be long to $\left\{|\psi_{m}\rangle\right\}$, another set of eigenstates $\left\{|\psi_{m}^{\prime}\rangle\right\}$ can be generated accordingly.
-Although $E_{0}$ can be real or complex, dependenting on the detail of the Hamiltonian $H_{0}$, we can conclude that the energy levels of Hamiltonian $H$ consists of multi-set of WSLs with an identical real level spacing $r_{0}\mathcal{\omega}$.
+Although $E_{0}$ can be real or complex, dependenting on the detail of the Hamiltonian $H_{0}$, we can conclude that the energy levels of Hamiltonian $H$ consists of multi-set of WSLs with an identical real level spacing $Nr\mathcal{\omega}$.
+In principle, this conclusion is true for any given $N$.
+However, a large energy level spacing that exceeds the scale of low-energy physics is meaningless.
+Fortunately, this conclusion drawn from few-body systems is applicable to specific invariant subspaces of many-body systems, in which energy eigenstates are localized.
+Here, we provide an example to demonstrate this point.
+Consider a system with a set of $n$-boson cluster eigenstates in the $n$-boson invariant subspace.
+These eigenstates are connected by the translational operator.
+In this situation, one could easily have a large number $K$ of widely separated $n$-boson clusters to construct the eigenstates with a total boson number $N=Kn$.
+The WSL frequency spacing derived above for states of that multi-cluster kind would therefore be $Knr\omega$.
+However, there exist many subsets of eigenstates, which can be constructed by translating one $n$-boson cluster while leaving the other $K-1$ clusters unchanged.
+Then, the corresponding energy levels in a subset form the WSL with frequency spacing $nr\omega$ rather than $Nr\omega$.
+This indicates that a system with a large boson number $N=Kn$ can also support dynamics with a smaller frequency $nr\omega$.
 Applying the conclusion to the several cases, the same result can be obtained in an exact manner.
@@
 (2016).
-This rigorous conclusion has important implications for investigating the dynamics of many-body and non-Hermitian systems in the presence of a linear field.
+This rigorous conclusion has important implications for investigating the dynamics of interacting and non-Hermitian systems in the presence of a linear field.
 In the following sections, we will focus on both the Hermitian and non-Hermitian extended Bose-Hubbard models to demonstrate the application of our conclusions and to reveal the dynamics of correlated bosons.
@@
 Parameters $\kappa$, $U$ and $V$, denote the tunneling strength, on-site and NN interactions between bosons.
-In this work, we consider the cases with real and imaginary , respectively.
+In this work, we consider the cases with real and imaginary $\kappa$, respectively.
 Obviously, Hamiltonian $H_{0}$ satisfies the conditions in the theorem, and then possesses multi-set of WSLs.
@@
 It is clear that, in the invariant subspace with fixed boson number $N$, there is a set of degenerate eigenstates $\left\{\left|l\right\rangle,l\in\left(-N\infty,N\infty\right)\right\}$ of $H_{0}\left(\kappa=0\right)$ with eigen energy $\varepsilon(N)$ $=UN(N-1)/2$, which are expressed as the form
-$\left|jn+\lambda\right\rangle=\frac{\left(b_{j}^{\dagger}\right)^{n-\lambda}\left(b_{j+1}^{\dagger}\right)^{\lambda}}{\sqrt{\left(n+1-\lambda\right)!\left(\lambda-1\right)!}}|\mathrm{vac}\rangle,$ (16)
+$\left|jN+\lambda\right\rangle=\frac{\left(b_{j}^{\dagger}\right)^{N-\lambda}\left(b_{j+1}^{\dagger}\right)^{\lambda}}{\sqrt{\left(N-\lambda\right)!\left(\lambda\right)!}}|\mathrm{vac}\rangle,$ (16)
 with $j\in\left(-\infty,\infty\right)$ and $\lambda\in[0,N-1]$.
@@
 This ensures us to perform the first-order perturbation approximation when the hopping term is considered under the strong correlation condition where $U\gg\left|\kappa\right|$.
-In order to obtain the effective Hamiltonian, one can simply envalue the nonzero matrix elements of $H_{\text{{B}}}$ in the subspace spanned by the set of eigenstates $\left\{\left|l\right\rangle\right\}$.
+In order to obtain the effective Hamiltonian, one can simply evaluate the nonzero matrix elements of $H_{\text{{B}}}$ in the subspace spanned by the set of eigenstates $\left\{\left|l\right\rangle\right\}$.
 For a given boson number $N$, the effective Hamiltonian can be written as
@@
 In this sense, the dynamics of $H_{\mathrm{eff}}^{[N]}$ is essentially the dynamics of $N$ correlated bosons.
+In Fig.
+1, a schematic illustration showing the correspondence between the configurations of the boson cluster with $N=3$ and the basis states for the effective Hamiltonian under strong interactions is presented.
 Notably, we would like to point out that the hopping strength in $H_{\mathrm{eff}}^{[N]}$ is of the order of $\kappa$, not $\kappa^{2}/U$.
@@
 The corresponding step for each ladder is $N\omega$.
+Figure 1: Schematic illustrations of the correspondence between the configurations of the boson cluster with $N=3$ and the basis states for the effective Hamiltonian given in Eq.
+17.
+(a1-a4) Four configurations of the boson cluster with $N=3$.
+These are degenerate states of the Hamiltonian $H_{0}(\kappa=0)$.
+Here, only the hopping processes within a dimer are presented.
+(b1-b4) The sites in the chain represent the basis states of the effective Hamiltonian.
+The hopping strengths indicated are non-uniform.
 Figure 2: Plots of $p_{l}(t)$ defined in Eq.
@@
 $H_{\mathrm{eff}}^{[4]}$ $=$ $-\kappa\sum_{j=-\infty}^{\infty}(2|4j\rangle\langle 4j+1|+\sqrt{6}|4j+1\rangle\langle 4j+2|\,$ (23)
-$+\sqrt{6}|4j+2\rangle\langle 4j+3|\,+2|4j+3\rangle\langle 4j+4|+\mathrm{H.c.})$
-$+\omega\sum_{l=-\infty}^{\infty}l\left|l\right\rangle\left\langle l\right|.$
+$+\sqrt{6}|4j+2\rangle\langle 4j+3|\,+2|4j+3\rangle\langle 4j+4|$
+$+\mathrm{H.c.})+\omega\sum_{l=-\infty}^{\infty}l\left|l\right\rangle\left\langle l\right|.$
 Numerical simulation shows that, when $\kappa$ is imaginary, there are two sets of real ladders and two sets of complex ladders.
-The two real ladders form a new ladder with energy levels given by ±$\pm 2\omega m$, where $m=0$, $1$, ….
+The two real ladders form a new ladder with energy levels given by ±$\pm 2m\omega$, where $m=0$, $1$, ….
 The real parts of two complex ladders is given by $(2\pm 4m)\omega$, where $m=0$, $1$, ….
@@
 V Dynamics of correlated bosons
-In this section, we focus on the investigation on the dynamics of correlated bosons.
-Considering the original Bose-Hubbard Hamiltonian as given by Eq.
-(12), the theorem regarding its energy levels applies to any set of parameters $\kappa$, $U$ and $V$.
-This feature can be characterized by the quantity $\varepsilon$, which is defined as the sum of the real part of $E$ and the imaginary part of $E$,
+According to the theorem, WSLs exist in a Bose-Hubbard model with arbitrary parameters $\kappa$, $U$ and $V$.
+However, the numerical results presented so far have been essentially for single-particle systems.
+In this section, we will focus on both the Hermitian and non-Hermitian extended Bose-Hubbard models with moderate values of $U$ and $V$.
+We will provide numerical results to show the existence of WSLs and present numerical results on the dynamics of correlated bosons, which offer evident demonstrations of our rigorous results.
+The feature of WSLs can be characterized by the quantity $\varepsilon$, which is defined as the sum of the real part of $E$ and the imaginary part of $E$,
 […]
@@
 In Fig.
-3, quantities $\varepsilon(\kappa)$ are plotted for both real and imaginary $\kappa$.
+4, quantities $\varepsilon(\kappa)$ are plotted for both real and imaginary $\kappa$.
 As expected, the plots of $\varepsilon(\kappa)$ are evidently several families of curves.
@@
 Such Hamiltonians consist of two distinct parts: an arbitrary Hamiltonian exhibiting translational symmetry and a linear potential.
-The most fascinating and important feature of these systems is that their conclusions are independent of the details of the Hamiltonian.
+The results are significant for few-body systems or for many-body systems with few-body clusters.
 As examples, we have investigate the extended Bose-Hubbard model with several bosons.
-Analytic and numerical approaches verified our conclusion in the case with $U=V$.
+Analytic and numerical approaches verified our conclusion in the cases with large and moderate $U=V$, real and imaginary $\kappa$.
 We also demonstrated the application of the theorem by analyzing dynamics driven by the Hubbard model using numerical simulations.
-Our findings provide a method for analyzing the energy-level structure of general systems and are expected to be both necessary and insightful for quantum engineering in many-body systems.
+Our findings provide a method for analyzing the energy-level structure of general systems and are expected to be both necessary and insightful for quantum engineering in quantum interacting systems.
 Acknowledgements.